If a production is Q=250K^0.7L^0.3 determine the marginal product of capital and marginal product of labour when k=50 and L=50

1
Expert's answer
2022-01-11T11:34:26-0500

"Q = 2 50K^{0.7} L^{0.3}"


The marginal product of labor = We will have to differentiate the above equation with respect to L


"\frac{dQ}{dL}= 0.7 × 250 K^{0.7}L^{0.3-1}"


 "\frac{dQ}{dL}= 0.7 × 250 K^{0.7}L^{-0.7}"


"\frac{dQ}{dL}=175\frac {K^{0.7 }}{L^{0.7}}"


"K=\ 50" and "L=\ 50"


Therefore:


Marginal Product of labor= 175.


"Q=250K^{0.7} L^{0.3}"

In order to find the Marginal product of capital, we have to differentiate the equation with respect to K.

Therefore:


"Q=250\times0.7K^{0.7-1} L^{0.3}"

"=\ 175K^{-0.3} L^{0.3}"

"=\ 175\frac{L^{0.3 }}{K^{0.3}}"


Whereas the value of K and L is the same as per the equation, therefore, the Marginal product of capital will be 175